Skip to Search

 
diagonal spacer

Information Technology Staff Meeting Notes
July 18, 2008

Present: Thomas Bielicke, Cynthia Bowker, Will Chaney, Logan Clapp, Matthew Daley, Debra Fetzer, Jason Fleming, Winston Harris, Paul McDonough, Bobbie Meng, Tom Minton, Patricia Ruwell, Mark Sullivan
Announcements
  1. Library IT staff meeting has been moved to the East conference room, Fridays 10AM-12PM.
  2. Bobbie has an air conditioning unit at her house now.
Review Minutes of 7/11/08
  1. Web #2: Name of the Linux browser is: Konqueror.
  2. Liaisons #3: Grover, not Glover.
Ongoing IT Updates
Equipment Received/Installed/Repaired
  1. Orders are flowing. Bill advised to 'order as normal'.
  2. Will had some questions about liaison/admin rights. By the end of the summer, they will no longer be able to keep their admin rights. There will be staff with special privileges. We'll need to make a list and document software that requires admin rights.
  3. Property IDs for the returned/replaced IBM laptops in MSL - Cynthia to email Bill Hanssen, giving him a list of the Property ID's that were returned to IBM and the SN of the new ones. We need to get replacement Property IDs for the new items received.
  4. Loaner laptop encryption - workstation support team to do it for loaners. Faculty laptops will need to be encrypted. Will have to query departments for different applications (such as ALEPH) that should be installed.
  5. CJK - almost done. Logan downloaded a demo copy and needs to purchase the license online.
Training
  1. Systems training - Grover training for staff on Monday and Tuesday next week. We will continue to train as needed.
  2. Workstation support unit - task lists, training, meetings. Evan will be out for two weeks. SWAT (Systems Workstation Action Team) will continue to meet and train.
Networking
  1. Disaster Prep - in progress.
  2. CNS - wall plate in room 310 has been defined to use either 54 or 223 subnet at port WA08 for Cathy Martyniak.
  3. CNS was in DLC - checking errors on the switches and verifying it is not a wiring problem.
  4. McAfee framework - done. Need to install Backup Exec Remote Agency on EPO.
  5. Greenstone backup - awaiting SAMBA configuration. Logan to check with Cynthia to see if it worked.
  6. DLCAV - moved SQL server there - it has files that need to be backed up.
  7. Maps server - backup? May need to buy tapes.
  8. DLC is discussing the purchase of a high-capacity tape drive. Logan is requesting a quote from IBM.
  9. SAN - corruption issue solved. VMware lost NTFS settings. Had to failover to the backup and rebuild the drives. Will fall back to the primary as soon as one more issue is solved.
  10. SAN migration - most of the smaller units have been migrated. We'll start with the bigger departments this week. Special Collections asked that their students have full access to their departmental folders. Mil made the request; we are awaiting response from the department chair.
  11. Campus AD - goal is to have public machines migrated by August 15. Details to follow.
  12. CNS wall plate edge protection rollout - on hold.
Software Projects Status
  1. Grover to go live for Systems on Monday, July 21 8AM. Everything will be double-entered until Tuesday at noon when it will be live for everyone. Will to take the Ticketmaster role in the event that both Debra and Cynthia are out.
  2. Tracking database - version in test - Erich to look at it.
  3. Greenstone - moved database to DLCAV. Failover between NT18 & DLCAV - SQL server Express functionality. Once the public machines are moved to AD, NT4 could go away.
  4. VIVO - Val and Sara sent the proposal to John and Judy. $3500 for hardware and to hire a programmer for a year. The hardware cost would be funded by the Library and we'd go to UF to fund the staff position.
  5. Betsy and Michelle Crump met with Winston to alleviate concerns about Jason and Michael moving to Library IT.
  6. Form-to-Mail programming change this week. We temporarily took it down. It's now back up.
  7. Discussion of slowness, corrupted files in Outlook2K7.
  8. Restart policy - forced reboots after updates.
  9. Shibboleth meeting today with Bridges.
  10. Bryant - Chris is working on it.
Web Projects
  1. IFAS station in Quincy changed their IPs not to use the Proxy server - now they can no longer access. Rich used to handle this: now who? Log the problem to Grover and it can go to both Jason and Rich.
Students and Department
  1. Xerox copier has been repaired: new tray, new feeder. No more noise.
  2. A new work study student, Shamika Brooks, is scheduled to start the fall semester on Aug. 19. That will be Debra's last day prior to vacation, so everyone will have to assist in helping to train her on the phones, office policy, etc.
  3. Rob will be here as OPS student through August 8. He may switch to OPS stadium for a week or two - to help out until after Labor Day.
IT Security Management
  1. Meeting held.

 

University of Florida Home Page